OpenStudy (anonymous):

anybody know the correct way to solve for Solve the system using the substitution method. y-3x=2 and y=(x+1)^2-5 I keep getting x= -1.688 and y is a number alike to that but i don't get a whole number

OpenStudy (mathstudent55):

Solve the left equation for y. What do you get? Then substitute that into the second equation.

OpenStudy (anonymous):

would i substitute (x+1)^2-5 - 3x = 2 for y in the first equation and solve?

OpenStudy (anonymous):

@mathstudent55 got it will do

OpenStudy (mathstudent55):

If you need more help, show what you get, and we can work on it together.

OpenStudy (anonymous):

y=3x+2 correct

OpenStudy (anonymous):

@mathstudent55 i got y =3x +2

OpenStudy (mathstudent55):

Good. Now substitute y in the second equation with 3x + 2. What do you get?

OpenStudy (anonymous):

@mathstudent55 set it up as 3x+2= (x+1)^2-5 right

OpenStudy (mathstudent55):

Great. Now you have one equation in one variable. We need to solve it for x. First square the binomial on the right sides.

OpenStudy (anonymous):

@mathstudent55 what is the binomial (x+1)^2-5?

OpenStudy (mathstudent55):

(x + 1)^2 is the binomial squared. You need to square it. Either do it by knowing the pattern of the squaring of a binomial \((a + b)^2 = a^2 + 2ab + b^2\), or use FOIL.

OpenStudy (anonymous):

@mathstudent55 x2+2x+1?

OpenStudy (aum):

Yes, (x+1)^2 = x^2 + 2x + 1 Put that in: 3x+2 = (x+1)^2- 5 and simplify

OpenStudy (anonymous):

@aum so i just simplify 3x+2 = (x+1)^2- 5

OpenStudy (aum):

3x+2 = (x+1)^2- 5 3x + 2 = x^2 + 2x + 1 - 5 3x + 2 = x^2 + 2x - 4 x^2 + 2x - 4 - 3x - 2 = 0 x^2 - x - 6 = 0 (x-3)(x+2) = 0 x = 3 or -2

OpenStudy (anonymous):

got it thanks @aum

OpenStudy (anonymous):

so how would i do this same equation using elimination method

OpenStudy (aum):

y = (x+1)^2 - 5 y - 3x = 2 Subtract the 2 equations and that would eliminate y: 3x = (x+1)^2 - 5 - 2 3x = (x+1)^2 - 7 (x+1)^2 - 7 - 3x = 0 x^2 + 2x + 1 - 3x - 7 = 0 x^2 - x - 6 = 0 (x-3)(x+2) = 0 x = -3, -2

OpenStudy (anonymous):

@aum so there is no solving for y in this one?

OpenStudy (aum):

You have to put each of the x value in either one of the equations and solve for y. Choose y - 3x = 2 which is the easier one.

OpenStudy (anonymous):

??

OpenStudy (anonymous):

so what you just typed Subtract the 2 equations and that would eliminate y: 3x = (x+1)^2 - 5 - 2 3x = (x+1)^2 - 7 (x+1)^2 - 7 - 3x = 0 x^2 + 2x + 1 - 3x - 7 = 0 x^2 - x - 6 = 0 (x-3)(x+2) = 0 x = -3, -2

OpenStudy (anonymous):

thats solving for x correct?

OpenStudy (aum):

Yeah that solves for x. But you are given two equations in two unknowns: x and y. You have solved for x and now you will have to find the corresponding y.

OpenStudy (anonymous):

@aum so how would i dove for y on the very first on using substitution ?

OpenStudy (aum):

Oh, I just noticed a typo in two of my earlier replies. In both cases: (x-3)(x+2) = 0 so x = 3; x = -2 (not -3).

OpenStudy (anonymous):

which ones was that 1. Subtract the 2 equations and that would eliminate y: 3x = (x+1)^2 - 5 - 2 3x = (x+1)^2 - 7 (x+1)^2 - 7 - 3x = 0 x^2 + 2x + 1 - 3x - 7 = 0 x^2 - x - 6 = 0 (x-3)(x+2) = 0 x = -3, -2 or 2. 3x+2 = (x+1)^2- 5 3x + 2 = x^2 + 2x + 1 - 5 3x + 2 = x^2 + 2x - 4 x^2 + 2x - 4 - 3x - 2 = 0 x^2 - x - 6 = 0 (x-3)(x+2) = 0 x = 3 or -2

OpenStudy (aum):

y-3x=2 or y = 3x + 2 put x = 3: y = 3(3) + 2 = 9 + 2 = 11 put x = -2: y = 3(-2) + 2 = -6 + 2 = -4 The solutions are: x = 3, y = 11; x = -2, y = -4 or (3,11) and (-2,-4) I guess the mistake was in one place only The one that has x = -3 should be x = 3.

OpenStudy (anonymous):

and that is for substitution or both ?

OpenStudy (anonymous):

@aum is that solving by substitution

OpenStudy (anonymous):

never mind i get it

OpenStudy (anonymous):

thank you i was confused now i know how to do this

OpenStudy (aum):

For both. Once you have solved for x (either by substitution method or the elimination method) you are done with those methods. After that you put x in the simpler of the two equations and compute y.

OpenStudy (anonymous):

do you know about logirithmics ?

OpenStudy (aum):

a little bit. This question has gotten too long. Close this one and post a new one.

OpenStudy (anonymous):

ok i will